Write True if the statement is correct and False if it is incorrect. Give explanation.

  1. Given H1: p ≠ 0.6, z = 0.78, then P-value = 2 P(z < 0.78).
  2. Given H1: p < 0.629, n = 19, and α = 0.05 then CV = -1.96.
  3. Given H1: μ < 2.8, n = 19, and α = 0.05 then CV = 1.734.
  4. If n = 696 and = 30%, then x = 209.
  5. If someone claims that the mean is over 523 and the test result rejects H0, then there is not sufficient evidence to support the claim that the mean is greater than 523.
  6. If the claim says that the mean IQ score of movie patrons is different from 100 and we reject H0 then we have enough evidence to reject the claim.
  7. If H1: µ < 14.7 and we reject H0 then there is not sufficient evidence to support the claim that the mean is less than 14.7.
  8. If n = 677 and x = 172, then = 0.26.
Answer #1

1) Given H1: p ≠ 0.6, z = 0.78, then P-value = 2 P(z < 0.78).

False.

P-value = 2*P(z > 0.78)

2) Given H1: p < 0.629, n = 19, and α = 0.05 then CV = -1.96.

False.

Cv = -1.645 for a = 0.05

3) Given H1: μ < 2.8, n = 19, and α = 0.05 then CV = 1.734.

True.

df = n-1 = 18

tdf,0.05 = t18,0.05 = 1.734

4) If n = 696 and = 30%, then x = 209.

True

p= x / n

x= np = 696 *0.30 =209

5) If someone claims that the mean is over 523 and the test result rejects H0, then there is not sufficient evidence to support the claim that the mean is greater than 523.

False.

If someone claims that the mean is over 523, Ho:- u = 523 & Ha:- u > 523

So we reject Ho means u > 523

there is sufficient evidence to support the claim that the mean is greater than 523.

6) If the claim says that the mean IQ score of movie patrons is different from 100 and we reject H0 then we have enough evidence to reject the claim.

False.

if we reject Ho then we have enough evidence to support the claim that the mean IQ score of movie patrons is different from 100

7) If H1: µ < 14.7 and we reject H0 then there is not sufficient evidence to support the claim that the mean is less than 14.7.

False.

we reject H0 then there is sufficient evidence to support the claim that the mean is less than 14.7.

8) If n = 677 and x = 172, then P= 0.26.

True

P= x / n = 172 / 677 = 0.2546 ~ 0.26